OpenStudy (anonymous):

a + 5c − b − c + a − b

OpenStudy (anonymous):

are you trying to simplify?

OpenStudy (anonymous):

yea

OpenStudy (marissalovescats):

Combine like terms :)

OpenStudy (marissalovescats):

a+a=? +5c-c=? -b-b=?

OpenStudy (anonymous):

okay 2a-2b+4c do you see how I got that? piggybacking off of what @marissalovescats said